40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 19 23:50:44 2013, Jenne, Update, LSC, PRCL locking again 
    Reply  Thu Jun 20 01:13:18 2013, Jenne, Update, LSC, PRCL locking again - no ASC success 
       Reply  Wed Jun 26 04:38:02 2013, Jenne, Update, LSC, PRCL locking again - ASC success PRMI_ASC_yawOnly_powerSpectra_25June2013.pdfPRMI_ASC_yawOnly_25June2013_mag.pdfPRMI_ASC_yawOnly_25June2013_phase.pdfPRMI_ASC_yawOnly_25June2013_coherence.pdfSensMatMeas_26June2013.png
          Reply  Thu Jun 27 17:09:41 2013, Jenne, Update, LSC, PRCL locking again - POP PIT work MeasuredVsModeledASCyaw.png
             Reply  Fri Jun 28 02:12:41 2013, Jenne, Update, LSC, PRCL locking again - ASC work PRMI_ASC_powerSpectra_27June2013.pdfScreenshot-Untitled_Window.png
Message ID: 8767     Entry time: Thu Jun 27 17:09:41 2013     In reply to: 8753     Reply to this: 8780
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRCL locking again - POP PIT work 

Last night before dinner, I copied over the ASC yaw servo filters to the ASC pitch filter bank.  Using ASC gain of +0.001, I was getting the ~250Hz oscillations that Rana and I had seen with yaw. 

Rana pointed out to me that my measured TF of the yaw loop doesn't look right up in the several hundred Hz region:

MeasuredVsModeledASCyaw.png

As you can see on the right side, which is all of the PRCL ASC yaw filter banks, multiplied by a simulated pendulum filter, the magnitude should just keep decreasing.  However, on the measured plot on the left, you can see that I have a little gain hump.  I'm not sure what this is from yet.

ELOG V3.1.3-